INFRARED SPECTROSCOPY AND MOLECULAR DYNAMICS IN OC-HI R. GARNICA, Z. WANG, A. I. MCINTOSH, R. R. LUCCHESE and \underline{J. W. BEVAN}, Department of Chemistry, Texas A&M University, College Station, Texas 77843.